Half of a number less than two is twenty.


1
Expert's answer
2021-10-04T19:24:39-0400

Let a number be "x"

Therefore, "\frac{x}{2}" "-2=20"

"\frac{x}{2}=20+2=22"

"x=22\times2"

"x=44"

Therefore, the number is 44.